a rectangular box is constructed to have the same volume as a cube whose edge is 6 cent. the rectangular box is 12 centimeters and the width is 8 centimeters what is the number centimeters in the height of the box
Answer by richwmiller(17219)

v=l*w*h
v=6*6*6
v=12*8*h
6*6*6=12*8*h
6*6*6=6*2*2*4*h
6*2*2*3*3=6*2*2*4*h
3*3=4*h
9=4h
2.5=height in centimeters
Reminder:
You are limited to one problem per submission.
No more than 4 submissions daily
No similar problems.